OpenStudy (anonymous):

The slope-intercept equation of a line is y = -7x - 2. What are the slope and y-intercept of the line? A. The slope is -7 and the y-intercept is (0, 2). B. The slope is -2 and the y-intercept is (0, -7). C. The slope is -7 and the y-intercept is (0, -2). D. The slope is -2 and the y-intercept is (0, 7).

OpenStudy (texaschic101):

In y = mx + b form, the slope can be found in the m position and the y intercept can be found in the b position. y = mx + b y = -7x - 2 slope(m) = -7 y int (b) = (0,-2)

OpenStudy (anonymous):

Thank you so much the way it was explained to me was horrible but this example helped

OpenStudy (texaschic101):

good to hear :)